OpenStudy (anonymous):

A man invests a certain amount of money at 2% interest and $800 more than that amount in another account at 4% interest. At the end of one year he earned $112 in interest. How much money was invested in each account? Which of the following equations could be used to solve the problem? 0.02x + 0.04(800) = 112 0.02x + 0.04(x + 800) = 112 0.04x + 0.02(x + 800) = 112

OpenStudy (tkhunny):

0.02x + 0.04(800) = 112 -- Bad. It's 800 MORE, not just 800. 0.02x + 0.04(x + 800) = 112 -- Looking possible. 0.04x + 0.02(x + 800) = 112 -- Backwards. The "More" should be at 4%.

OpenStudy (anonymous):

so which one is it the third one

OpenStudy (anonymous):

yes.

OpenStudy (tkhunny):

What?? That makes so much no sense. Look at the discussion. Understand the comments. The first and third are discounted, disparaged, and thrown out. The middle one is the only remaining possibility.

OpenStudy (anonymous):

im confused im getting 2 different responses

OpenStudy (anonymous):

If you work it out then the only one that works is the third one.

OpenStudy (tkhunny):

{sigh} The third one is correct if you don't mind getting the rates backwards. Read my first post very carefully and make up your own mind. You needn't be confused.